OpenStudy (anonymous):

(x)/(x-4)-(4)/(x-4)=(x^2+16)/(x^2-16) x = ?

OpenStudy (dumbcow):

Combine fractions on left side \[\frac{x-4}{x-4} = \frac{x^{2}+16}{x^{2}-16}\] Simplify \[1 = \frac{x^{2}+16}{x^{2}-16}\] \[x^{2} -16 = x^{2} +16\] \[-16 = 16\] This is False, thus x has no solution